# Content-Style Disentangled Representation for Controllable Artistic Image Stylization and Generation

### v1

@@ Abstract
Controllable artistic image stylization and generation aims to render the content provided by text or image with the learned artistic style, where content and style decoupling is the key to achieve satisfactory results. However, current methods for content and style disentanglement primarily rely on image information for supervision, which leads to two problems: 1) models can only support one modality for style or content input;2) incomplete disentanglement resulting in semantic interference from the reference image. To address the above issues, this paper proposes a content-style representation disentangling method for controllable artistic image stylization and generation. We construct a WikiStyle+ dataset consists of artworks with corresponding textual descriptions for style and content. Based on the multimodal dataset, we propose a disentangled content and style representations guided diffusion model. The disentangled representations are first learned by Q-Formers and then injected into a pre-trained diffusion model using learnable multi-step cross-attention layers for better controllable stylization. This approach allows model to accommodate inputs from different modalities. Experimental results show that our method achieves a thorough disentanglement of content and style in reference images under multimodal supervision, thereby enabling a harmonious integration of content and style in the generated outputs, successfully producing style-consistent and expressive stylized images. The code of our method will be available upon acceptance.
@@ 1 Introduction
Artistic image stylization and generation task aims at creating new images by applying specific artistic styles to content. In stylization, the concept of “content” is well-defined, typically referring to the subject of an image or the semantics of input text. However, the definition of “style” is relatively vague and lacks consistent standards. Art historian, Meyer Schapiro, has defined artistic style as: “The constant form and sometimes the constant elements, qualities, and expression in the art of an individual or a group” (Karkov & Brown, 2003). For example, Impressionism emphasizes the natural representation of light and color. Therefore, to truly capture and reproduce artistic style, stylization models must learn the unique ways of artists handle and present content in their creative process. To achieve this, the model needs to effectively disentangle and control content and style during the stylization process, thereby producing visually coherent and stylized results.
Recently, diffusion models (Ho et al., 2020; Rombach et al., 2022) have demonstrated great potential in text-to-image stylization tasks (Ramesh et al., 2022; Ye et al., 2023; Mou et al., 2024; Chen et al., 2024a) due to their powerful generative capabilities. These methods typically extract reliable features in the reference images serving as conditional information to guide the diffusion model to follow the predetermined style. However, the features extracted by the encoder often couple style and semantics, leading to semantic conflicts between the text and the reference image . This in turn causes content leakage from the reference image, as shown in Fig. 1 (a). Some approaches (Xing et al., 2024) (Qi et al., 2024) attempt to achieve disentanglement by using separate encoders to extract style and content representations. Still, they encounter the following issues: 1) These methods depend on datasets consisting of paired content and stylized images for training. Besides, the disentanglement process is solely reliant on image-based supervision, which limits the style input to only image modality. 2) They primarily focus on extracting style representations from the reference image but do not effectively decouple the content and style. This results in a mismatch between style and content in the generated images, failing to achieve the desired stylization effect (as shown in Fig.1 1(b)). Additionally, they can only handle visual elements from single image and unable to mimic the style of an artist or genre. Artbank (Zhang et al., 2024) achieves artist-based stylization by constructing a style prompt bank to store knowledge from art collections, which needs to be retrained when extend to new artists. However, Artbank is trained with the reconstruction objective with image-based supervision, which leads to unsatisfied performance in text-controlled image generation.
@@
To address the aforementioned issues, this paper proposes a content-style representation disentanglement method for controllable artistic image stylization and generation. First, we constructed a multimodal art image-text dataset, WikiStyle+. We collected art images and their corresponding style information from the WikiArt website, including details such as artists, creation periods, genres, and painting medias, as references for style descriptions. Additionally, we used large language model to generate textual descriptions of the content of these artworks. In this way, we tackled the problem of the lack of explicitly disentangled data for style and content in artistic images from a multimodal perspective. Based on the constructed WikiStyle+ dataset, we proposed a disentangled content and style representations guided diffusion model. Through multimodal alignment tasks, the Q-former aligns the learned image style features with style descriptions and the learned content features with content descriptions. This method utilizes multimodal data to provide the style and content information of the reference images for disentanglement supervision, achieving explicit separation of content and style. As a result, the model learns style representations consistent with the definition of style in art history, which reflects the distinct ways artists depicts the world. Building on this, we inject the learned style and content representations into the multi-step cross-attention layers of the diffusion model, enabling text-to-image stylization and generation.
With the explicit disentanglement of style and content, our method can adjust visual elements based on content and different style prompts, making stylization go beyond mere visual imitation and achieve a harmonious integration of content and style in the generated results, as shown in Fig. 1 (c). Our approach can effectively prevent content leakage from reference image. Besides, the multimodal disentangle supervision allows us to accommodate style and inputs from either image or text modalities flexibly, which is typically restricted by previous image-supervised decoupling methods (e.g., DeaDiff Qi et al. (2024)). In summary, our contributions are threefold:
@@
We constructed the art image-text dataset, WikiStyle+, addressing the problem of the lack of explicitly disentangled data for style and content from a multimodal perspective.
@@
We propose a disentanglement method that explicitly decouples style and content representations through multimodal supervision, enabling the model to accept inputs from different modalities as control conditions.
@@
We constructed a disentangled representation-guided diffusion model, where the disentangled content and style representations are injected into the cross-attention layers at different time steps of the diffusion model, successfully generating images with harmonious integration of content and style.
@@ 2.1 Diffusion-based image stylization with multimodal latents
In recent years, Diffusion Probabilistic Models Sohl-Dickstein et al. (2015) have demonstrated immense potential in the field of image generation Dhariwal & Nichol (2021); Ho et al. (2020); Song et al. (2020).Thanks to large-scale multimodal pre-trained models Radford et al. (2021); Li et al. (2023), diffusion models have achieved remarkable success in text-to-image generation tasks Ramesh et al. (2022); Rombach et al. (2022); Saharia et al. (2022). Text-to-image stylization tasks based on diffusion models are primarily divided into two categories: one is optimization-based methods, which involves fine-tuning the diffusion model to enable it to generate images for a specific style Ruiz et al. (2023); Kumari et al. (2023); Sohn et al. (2023), or utilizing textual inversion to redefine or optimize textual embeddings related to a specific style Gal et al. (2022); Zhang et al. (2023), thereby generating more accurately stylized images that match with descriptions. The other category involves conditional diffusion models based on pre-trained encodersHuang et al. (2023); Li et al. (2024); Mou et al. (2024); Wang et al. (2023b); Ye et al. (2023); Zhao et al. (2024), which extract stylistic features from reference images using pre-trained encoders and then inject these features as conditions into diffusion model to generate images with specific styles.
2.2 Content-Style Decoupling
In image stylization tasks, decoupling content and style is key to enhancing the quality of generated images. Text inversion methods (such as InST (Zhang et al., 2023), VCT (Cheng et al., 2023), DreamBooth (Ruiz et al., 2023), Textual Inversion (Gal et al., 2022), Custom Diffusion (Kumari et al., 2023), and ArtBank (Zhang et al., 2024)) map reference images the embedding space of special text tokens through a reversal module. They construct text prompts to specify whether the special tokens represent content or style, thereby separately decoupling content and style, such as “a dog in [v] style” or “a [v] dog in Van Gogh style,” where [v] represents either style or content information.
For unoptimized attention mechanism methods like StyleAligned (Hertz et al., 2024) and Visual Style Prompting (Jeong et al., 2024) typically achieve zero-shot content-style decoupling for image stylization by sharing or optimizing attention mechanisms in SD. Cross-domain alignment methods like StyleDiffusion (Wang et al., 2023a) and OSASIS (Cho et al., 2024) use cross-domain and intra-domain losses in CLIP space to achieve latent content-style decoupling.
Adapter-based methods (such as IP-Adapter (Ye et al., 2023), T2I-Adapter (Mou et al., 2024), and StyleAdapter (Wang et al., 2023b)) adjust specific layers or channels of the model, allowing it to effectively separate and blend style and content without altering its original architecture. Adapter methods based on pre-trained encoders (such as InstantStyle (Wang et al., 2024) and DEA-Diffusion (Qi et al., 2024)) decouple the content and style of reference images in the encoding space. However, current methods fail to achieve a thorough decoupling of content and style from an artistic creation perspective, resulting in content leakage from the reference image in the generated results.
3 Dataset
3.1 Dataset Construct Pipeline
A significant challenge in content-style decoupling is the lack of paired dataset. Current methods construct content-stylized image pairs using proposed prompts with pretrained text-to-image generation models. However, due to the randomness of generative models, even when using the same prompt, there are significant differences between the style and content of the images generated, necessitating tedious manual data cleaning. Fortunately, existing large language models have demonstrated a strong fidelity in generating image descriptions. We can construct a content description-style description-artwork triplet dataset to addressing the problem of the lack of explicitly disentangled data for style and content from the multimodal perspective. Moreover, the impact of randomness in text generation on semantics is far less than the impact of randomness in image generation on style.
Formally, the construction of the paired datasets involves the following three steps:
Step 1. Art image collection. We collect 189,631 data entries from the WikiArt websiteWikipedia (2021). The gathered data includes artistic images, along with detailed professional information provided by WikiArt for each artwork. This information pertains style-related attributes like the author, style, genre, media, and other relevant labels.
Step 2. Art image selection.Upon analyzing the data, we observe that works like photography, architecture, design drawings, advertisements, and illustrations often lack distinct stylistic features, while art movements like abstract and minimalism usually lack specific subjects. Based on these findings, we filter out works labeled under these genres, ultimately forming an art dataset comprising 146,547 records.
@@ 3.1 Dataset Construct Pipeline
Step 3. Content description generation.To construct paired content and style data, we utilized the recently open-sourced InternVL-Chat Chen et al. (2024b) with the prompt, “<image>, describe the content of this picture briefly.” to generate content descriptions for each artwork. This approach enabled us to obtain both the content and style information of the artistic images from a multimodal perspective, effectively addressing the issue of insufficient explicit data for content-style disentanglement.
3.2 Dataset Details
The final dataset comprises 146,547 triplets of art images, style descriptions, and content descriptions. Based on the professional annotation on WikiArt, we selected 4 style attributes under different categories, including artist (2,789), artistic style (209), genre (63), and mediums (184). In terms of content, it covers a variety of themes such as portraits, still lifes, natural landscapes, and cultural landscapes. More details about WikiStyle+ are provided in Appendix A.1.
@@ 4.1 Content and Style Disentanglement Network
The dataset is structured as triplets (image, style text, content text) as introduced in Sec. 3, which gives us data basis for disentangling content and style altogether. The style text is created by concatenating fields related to style in the dataset, such as artist, date, style and so on, in the format of "<key >is <value >". The content text is derived from the description field in the dataset, generated by a multimodal large language model. We achieve content and style disentanglement by minimizing an objective function that incorporates the Image-Text Contrastive Learning Loss , Image-Text Matching Loss , and Image-grounded Text Generation Loss as follows:
@@
Remark: To prevent the model from becoming complacent during training by simply copying content from image features, which could diminish its generalization ability, we randomly replace the style or content embeddings extracted from images with the corresponding embeddings extracted from texts. Both content and style embeddings are randomly selected from multiple modalities to participate in training the model. The logic behind this approach lies in feeding more complex tasks to train the model is beneficial for improving model’s capabilities. This approach fosters a more nuanced and versatile understanding of both content and style, enabling the model to generate images that are both faithful to the input content and creatively infused with the desired style. Additionally, it also enables our model to generate outputs with various multimodal combinations. Finally, we randomly drop the keywords of style texts for enabling the model to accept a wider variety of style keyword combinations when using text as style prompts.
@@ 5.1 Experiment Settings
Implementation Details. We trained on L20-40G GPUs with 3750 total batches, using AdamW (Loshchilov, 2017) as the optimizer, with a learning rate of 5e-5, and performed 100 iterations. Regarding reasoning, for guidance without classifiers (Ho & Salimans, 2022), we use a scale of 7.5 and set T = 50 steps for DDIM (Song et al., 2020) sampling. All comparison methods were implemented using publicly available code and default settings. Detailed settings of our model are elaborated in Sec. A.2.
Evaluation Metrics. Based on previous works (Qi et al., 2024) (Jiang & Chen, 2024), the following metrics are used in the style transfer task. For content similarity, we use text alignment capability (TA), which calculate the cosine similarity in the CLIP embedding space between stylized images and their corresponding text prompts. For Image Quality (IQ), the LAION-Aesthetics Predictor V2 1 is adopted to evaluate the generation quality of the image. For style similarity (SS), we use a text prompt template to express style, namely “the painter is [v], the theme is [v]”. Then, we measured the cosine similarity between the generated image and the corresponding style template in the CLIP text-image embedding space. We also conduct a user study to reflect the subjective preference (SP) for the results.
@@ 5.2 Comparison with State-of-the-Arts
In this section, we compare our method with the state-of-the-art methods. We introduce the experimental results based on supported input modalities and tasks as follows: 1) Text-to-image stylization, including optimization-based methods: InST (Zhang et al., 2023), encoder-based methods: IP-Adapter (Ye et al., 2023), DeaDiff (Qi et al., 2024), T2I-Adapter (Mou et al., 2024), and InstantStyle (Wang et al., 2024); 2) Text-to-image Generation, featuring advanced models DallE (Ramesh et al., 2021) and Stable Diffusion (SD) (Rombach et al., 2022); 3) Collection-based stylization: Artbank (Zhang et al., 2024). More results please refer to Sec. A.3.
5.2.1 Text-to-image stylization
Fig.4 illustrates the comparison results with state-of-the-art methods. For methods lacking effective decoupling mechanisms, such as IP-Adapter, T2I-Adapter and InST, semantic conflicts from the reference images are evident in the generated results, as shown in the first and third rows of Fig. 4. DEADiff and InstantStyle struggle to accurately mimic the style of the reference image when there is significant semantic gap between the reference image and the textual prompt. In contrast, our method not only faithfully follows textual prompts but also significantly mimics the artistic style contained in the reference images, making the stylized images closer to the true creations of the artists.
5.2.2 Stylized Text-to-image Generation
Fig. 5 compares our method with advanced text-to-image generation methods. Due to the lack of a decoupling mechanism, Dall-E’s generated images exhibit repetitive style patterns. Furthermore, the textures in the generated images do not resemble oil paintings. Our method, when prompted with different artists, generates distinct styles and adheres closely to the textual prompts, achieving performance comparable to that of Stable Diffusion. Due to our method’s use of multimodal supervision during the pre-training phase, it also performs well in text-to-image generation tasks.
5.2.3 Collection-based stylization
Fig. 6 presents a visual comparison between our method and ArtBank. Our approach shares a common feature with Artbank in that it learns the style information of art collections rather than the style of individual images. From Fig. 6, it is evident that our outputs more closely align with the artist’s style and the content text. For instance, in first row, our results more closely resemble Van Gogh’s portrait style compared to Artbank. In the stylized results with subject “barbecue”, Artbank’s outputs exhibit content leakage from the reference image, whereas our results effectively capture the semantic of content prompt. Since Artbank also employs a reconstructive training approach, the image structure tends to be unclear when text is used as the content prompt, as demonstrated in the second and third rows of the images.

### v2

# WikiStyle+: A Multimodal Approach to Content-Style Representation Disentanglement for Artistic Image Stylization

@@ Abstract
Artistic image stylization aims to render the content provided by text or image with the target style, where content and style decoupling is the key to achieve satisfactory results. However, current methods for content and style disentanglement primarily rely on image supervision, which leads to two problems: 1) models can only support one modality for style or content input;2) incomplete disentanglement resulting in content leakage from the reference image. To address the above issues, this paper proposes a multimodal approach to content-style disentanglement for artistic image stylization. We construct a WikiStyle+ dataset consists of artworks with corresponding textual descriptions for style and content. Based on the multimodal dataset, we propose a disentangled representations-guided diffusion model. The disentangled representations are first learned by Q-Formers and then injected into a pre-trained diffusion model using learnable multi-step cross-attention layers. Experimental results show that our method achieves a thorough disentanglement of content and style in reference images under multimodal supervision, thereby enabling more refined stylization that aligns with the artistic characteristics of the reference style. The code of our method will be available upon acceptance.
@@ 1 Introduction
Artistic image stylization task aims at creating new images by applying artistic styles to content. In stylization, the concept of “content” is well-defined, typically referring to the subject/semantics of the input image or text. However, the definition of “style” is relatively vague and lacks consistent standards. Art historian, Meyer Schapiro, has defined artistic style as: “The constant form and sometimes the constant elements, qualities, and expression in the art of an individual or a group” (Karkov and Brown, 2003). For example, Impressionism emphasizes the natural representation of light and color. Therefore, to truly capture the artistic characteristics of the reference style, stylization models needs to effectively disentangle and control content and style representation during the stylization process.
Recently, diffusion models (Ho et al., 2020; Rombach et al., 2022) have demonstrated great potential in text-to-image stylization tasks (Ramesh et al., 2022; Ye et al., 2023; Mou et al., 2024; Chen et al., 2024a) with their powerful generative capabilities. These methods typically extract reliable features in the reference images serving as conditional information to guide the diffusion model to follow the predetermined style. However, the features extracted by the encoder often couple style and semantics. This, in turn, results in content leakage within the stylized output, where elements from the reference image appear in the generated result despite being inconsistent with the intended subject, as shown in Fig. 1 (a).
Some approaches (Xing et al., 2024) (Qi et al., 2024) attempt to achieve disentanglement by using separate encoders to extract style and content representations. However, they still face the following challenges: 1) These methods depend on AI-generated paired content and stylized images rather than artistic works, which in turn causes the generated results resemble cartoon-like visuals rather than authentic oil paintings (Fig.1(b)). Moreover, they rely solely on image-based supervision, limiting the model to image inputs. 2) These methods overlook the fact that artists adapt their style based on different subjects, leading to generated images that exhibit similar colors and textures regardless of the theme (Fig.1(b)). This inconsistency fails to reflect the way real artworks vary in style across different subjects.
@@
To overcome these challenges, we introduces a multimodal approach for content-style representation disentanglement. First, we constructed a multimodal artistic dataset, WikiStyle+. We curated authentic art images from the WikiArt website along with their associated style information, including artist names, genres, and painting mediums, to serve as references for style descriptions. Additionally, we utilized a large language model to generate textual descriptions of the content depicted in these artworks. In this way, we addressed the lack of explicitly disentangled style and content data in artistic images from a multimodal perspective. Based on the constructed WikiStyle+ dataset, we proposed a content-style disentangled representation-guided diffusion model. Through multimodal alignment tasks, the Q-former aligns the learned image style features with style descriptions and the learned content features with content descriptions. This method utilizes multimodal data to provide the style and content information of the reference images for disentanglement supervision, achieving explicit separation of content and style information from art images. Building on this, we inject the learned style and content representations into the multi-step cross-attention layers of the diffusion model, leveraging its generative capabilities to achieve image stylization and generation.
With the explicit disentanglement of style and content, our method adapts visual elements based on the subject and varying style prompts, moving beyond simply replicating the color palette of the reference image. This enables more nuanced stylization that better captures the artistic characteristics of the intended style. As shown in Fig. 1 (c), our method captures broader brushstroke techniques of the artist rather than simply replicating signature swirling patterns. Moreover, the multimodal disentanglement supervision enables style and content inputs from both image and text modalities, overcoming the limitations of previous methods that rely solely on image-based supervision for decoupling. In summary, our contributions are threefold:
@@
We constructed WikiStyle+, a multimodal artistic dataset, to address the lack of explicitly disentangled style and content data from a multimodal perspective.
@@
We propose a multimodal approach for the explicit disentanglement of style and content representations through multimodal supervision, enabling the model to accept diverse modality inputs for artistic image stylization.
@@
We proposed a disentangled representation-guided diffusion model, where the disentangled content and style representations are injected into the cross-attention layers at different time steps of the diffusion model, enabling more refined stylization that aligns with the artistic characteristics of the reference style.
@@ 2.1 Diffusion-based image stylization with multimodal latents
In recent years, Diffusion Probabilistic Models Sohl-Dickstein et al. (2015) have shown great potential in image generation Dhariwal and Nichol (2021); Ho et al. (2020); Song et al. (2020). With the advancement of large-scale multimodal pre-trained models Radford et al. (2021); Li et al. (2023), diffusion models have achieved remarkable success in text-to-image generation Ramesh et al. (2022); Rombach et al. (2022); Saharia et al. (2022). Text-to-image stylization using diffusion models primarily falls into two categories: Optimization-based methods, which fine-tune diffusion models for specific styles Ruiz et al. (2023); Kumari et al. (2023); Sohn et al. (2023) or use textual inversion to refine text embeddings for improved style fidelity Gal et al. (2022); Zhang et al. (2023). Conditional diffusion models with pre-trained encoders, which extract stylistic features from reference images Huang et al. (2023); Li et al. (2024); Mou et al. (2024); Wang et al. (2023a); Ye et al. (2023); Zhao et al. (2024) and inject them as conditions into the diffusion model to generate images in the desired style.
2.2 Content-Style Disentanglement
In image stylization, content and style disentanglement is crucial for accurately transferring the target style. Text inversion methods, such as InST (Zhang et al., 2023), VCT (Cheng et al., 2023), DreamBooth (Ruiz et al., 2023), and ArtBank (Zhang et al., 2024), map reference images to the embedding space of special text tokens via a reversal module. They construct text prompts to separately provide content and style. Unoptimized attention-based methods, like StyleAligned (Hertz et al., 2024) and Visual Style Prompting (Jeong et al., 2024), achieve zero-shot content-style separation by modifying attention mechanisms in Stable Diffusion (SD). Cross-domain alignment approaches, such as StyleDiffusion (Wang et al., 2023b) and OSASIS (Cho et al., 2024), leverage cross-domain and intra-domain losses in CLIP space to decouple style and content at a latent level. Adapter-based methods, including IP-Adapter (Ye et al., 2023) and StyleAdapter (Wang et al., 2023a), fine-tune specific model layers or channels to separate and blend style-content without altering the model’s core structure. Similarly, encoder-based adapters, like InstantStyle (Wang et al., 2024) and DEA-Diffusion (Qi et al., 2024), disentangle style-content within the encoding space. Despite these advances, existing methods fail to achieve explicit disentanglement in latent space, leading to content leakage, where unintended elements from the reference image persist in the generated output.
3 WikiStyle+ Dataset
A major challenge in content-style disentanglement is the lack of proper supervision. Existing methods rely on AI-generated paired content and stylized images rather than genuine artistic works. However, real artworks typically do not have corresponding content images paired with them, making supervised disentanglement more difficult. In this paper, we propose a "content description"–"style description"–"artwork" triplet dataset to address the lack of explicitly disentangled style and content data from a multimodal perspective. Formally, the construction of the paired datasets involves the following three steps:
Step 1. Image collection. We collected 189,631 entries from WikiArt Wikipedia (2021), encompassing diverse artists and art movements. Each entry includes an artwork image and style-related metadata such as the artist, style, genre, medium, and other relevant attributes.
Step 2. Dataset Refinement. To ensure stylistic clarity, we removed artworks lacking distinct stylistic features, such as those in photography, architecture, design drawings, and advertisements. Additionally, we excluded works without clear subjects, including those from abstract and minimalism movements. After this filtering process, we obtained a refined dataset of 146,547 records. For style attributes, we selected the four most prevalent and distinctive ones: artist, artistic style, genre, and medium.
@@ 3 WikiStyle+ Dataset
Step 3. Content description generation. To construct paired content and style data, we utilized InternVL-Chat Chen et al. (2024b) with the prompt, “<image>, describe the content of this picture briefly." to generate content descriptions for each artwork. The generated content spans a diverse range of themes, including portraits, still lifes, natural landscapes, and cultural landscapes.
In this way, we construct a content-style dataset based on real artistic images from a multimodal perspective, addressing the issue of insufficient explicit data for content-style disentanglement. Further details on WikiStyle+ are provided in Supplementary Material Sec.1.
@@ 4.1 Content and Style Disentanglement Network
The dataset is structured as triplets as introduced in Sec. 3, which gives us data basis for disentangling content and style altogether. We achieve content and style disentanglement by minimizing an objective function that incorporates the Image-Text Contrastive Learning Loss , Image-Text Matching Loss , and Image-grounded Text Generation Loss as follows:
@@
Remark: To prevent the model from becoming complacent during training by simply copying content from image features, which could diminish its generalization ability, we randomly replace the style or content embeddings extracted from images with the corresponding embeddings extracted from texts. Both content and style embeddings are randomly selected from multiple modalities to participate in training the model. The logic behind this approach lies in feeding more complex tasks to train the model is beneficial for improving model’s capabilities.
This approach fosters a more nuanced and versatile understanding of both content and style, enabling the model to generate images that are both faithful to the input content and creatively infused with the desired style. Additionally, it also enables our model to generate outputs with various multimodal combinations. Finally, we randomly drop the keywords of style texts for enabling the model to accept a wider variety of style keyword combinations when using text as style prompts.
@@ 5.1 Experiment Settings
Implementation Details. We trained on L20-40G GPUs with 3750 total batches, using AdamW Loshchilov (2017) as the optimizer, with a learning rate of 5e-5, and performed 100 iterations. Regarding reasoning, for guidance without classifiers (Ho and Salimans, 2022), we use a scale of 7.5 and set T = 50 steps for DDIM (Song et al., 2020) sampling. All comparison methods were implemented using publicly available code and default settings. Detailed settings of our model are elaborated in Supplementary Material Sec.2.
Evaluation Metrics
Following previous works (Qi et al., 2024; Jiang and Chen, 2024), we evaluate multi-modal stylization using the following metrics:
Content Similarity (TA): Measured by the cosine similarity between stylized images and their corresponding text prompts in the CLIP embedding space. Image Quality (IQ): Assessed using the LAION-Aesthetics Predictor V2 1. Style Similarity (SS): Defined by a text prompt template "the painter is [v], the theme is [v]", with similarity computed between the generated image and the style prompt in the CLIP text-image embedding space. Subjective Preference (SP): Evaluated through a user study.
@@ 5.2 Comparison with State-of-the-Arts
In this section, we compare our method with the state-of-the-art methods. We introduce the experimental results based on supported input modalities and tasks as follows: 1) Text-to-image stylization, including optimization-based methods: InST (Zhang et al., 2023), encoder-based methods: IP-Adapter (Ye et al., 2023), DeaDiff (Qi et al., 2024), T2I-Adapter (Mou et al., 2024), and InstantStyle (Wang et al., 2024); 2) Text-to-image generation, featuring advanced models DallE (Ramesh et al., 2021) and Stable Diffusion (SD) (Rombach et al., 2022); 3) Collection-based stylization: Artbank (Zhang et al., 2024); 4) Content-style disentanglement: DEADiff (Qi et al., 2024), StyleDrop Sohn et al. (2023) and DreamStyler Ahn et al. (2024) Due to space limitations, the comparisons for 2-4 are provided in Supplementary Material Sec.5.1-5.3.
@@
5.2.1 Text-to-image stylization
Fig.4 illustrates the comparison results with state-of-the-art methods. For methods lacking effective decoupling mechanisms, such as IP-Adapter, T2I-Adapter and InST, semantic conflicts from the reference images are evident in the generated results, as shown in the first and third rows of Fig. 4. DEADiff and InstantStyle struggle to accurately capture the style of the reference image when there is significant semantic gap between the reference image and the textual prompt. In contrast, our method not only faithfully reflects textual prompts but also aligns with the artistic characteristics of the reference style. For more comparison results, please refer to Supplementary Material Sec 5.4.